Find the equation of the cone with the vertex at (1,−1,2) and the base curve as
(z+1)^2 = x+2, y = 3.
1
Expert's answer
2020-02-18T03:36:40-0500

Answer: "6y^2+24zy-4yx-32y+16z^2+26-40z-4x"


The vertex is "A(1, -1, 2)" . Let "a, b, c" be the direction ratios of a generator of the cone. Then the equations of generator are,


"\frac{x-1}{a}=\frac{y+1}{b}=\frac{z-2}{c}=t" (say)


The coordinates of any point on the generator are ("1+at" , "-1+bt" , "2+ct"). For some "t \in \mathbb{R}" , ("1+at" , "-1+bt" , "2+ct" ) lies on the guiding curve.

Therefore "((2+ct)+1)^2 = (1+at)+2" and "-1+bt = 3" . Thus, "t = \frac{4}{b}" . From this we get,


"((2+c\cdot\frac{4}{b})+1)^2 = (1+a\cdot\frac{4}{b})+2"


"(3+4\cdot\frac{z-2}{y+1})^2 = 3+4\cdot\frac{x-1}{y+1}"


After simplification, the required equation of the cone is,

"6y^2+24zy-4yx-32y+16z^2+26-40z-4x".